LHC 2011 - latest news

Week 42: Machine coordinators: B. Holzer & M. Lamont

Goal of the week: Luminosity production, MDs (90 m), 24 hours of 25 ns penciled in for Monday 24th October




Link to report from morning meeting

Monday 24th October

  • fill 2242 dumped around 5:00 Monday - trip of RCBX.V2.R1(again)
  • prepare for 25 ns tests (24 hours)

08:30 meeting

Special running
Massi Ferro-Luzzi

First look at TOTEM results Karsten Eggert

High bunch intensity instability Elias Metral


Sunday 23th October

  • 01:45h dump: failure of RCBX.V2.R1
  • 03:56h injection
  • 05:45h luminosity for a couple of hours, fill lost to QPS problem Q9
  • bit of a struggle with access system Sunday morning
  • 18:00 ramp > 2e14, into stable beam without any problems, 3.5e33 cm-2s-1
09:00 meeting

Saturday 22th October

  • Overnight Friday to Saturday: loss maps, injection issues, 06:45 Stable beams
  • 07:45 Loss to Stable beam flag glitch
  • 08:34h injection, ramp, squeeze
  • 11:14h luminosity
09:00 meeting

Friday 21th October

  • 02:20h injection, ramp un-squeeze
  • qps trigger in s45, tune trim quads
  • new attempt, successfull, van der meer scans until early afternoon - beams lost to trip of inner triplet power converter L2.
  • 25 ns run postponed until Monday
  • 17:40 Electrical network glitch - 400 kV manipulation by EDF - Static Var Compensators down around ring, lose Alice, LHC dipoles and ATLAS's torroid.
  • 22:00 Back in business
08:30 meeting

Thursday 20th October

  • Switch to Van der Meer scans overnight - unsuccessfully, beams lost in squeeze to software interlock on orbit corrector strength
  • Back to TOTEM/ALFA mode - back through unsqueeze by 09:00 Thursday morning !
  • 11:07h pots at 6.5 sigma, data taking
  • 18:00h vertical pots approached to 5.5 sigma data taking.
  • 22:00h fill end, dump, precycle
  • 22:22h power converter trip, access, change of pc
08:30 meeting

Wednesday 19th October

  • Lower intensity beams unsquoozen overnight for optics measurements at 90 m
  • Refill, ramp, unsqueeze for TOTEM, ALFA data taking
  • Beam dumped at 16:40 - self trigger of beam dump
  • Back in unsquoozen collisions by 20:35
  • Beams lost just at start of data taking - power converter L4, access required.
08:30 meeting

Tuesday 18th October

  • 02:10 Fill lost to brief loss of cryo maintain due to a faulty temperature reading caused a slow power abort on the main circuits in S81
  • Injection kickers a bit warm back at 450 GeV - 4 hours to cool
  • Loss maps at 450 GeV
  • 90 m setup during the day, some data taking but high background - bad beam quality, finished with loss maps
08:30 meeting

Monday 17th October

  • 07:00 Dump fil #2219, 123 pb-1 delivered.
  • Access for QPS
  • Power converter tests at 450 GeV
  • 14:11 Stable beams 3.6e33, 1.41e11 ppb
  • 15:45 Beam dump - power converter fault
  • Next fill lost in adjust, up to 1.45 ppb - looked like head-tail instablity beam 2
  • Back into stable beams by 22:10 with slightly reduced bunch current
08:30 meeting